# Slaves to the Law of Large Numbers: An Asymptotic Equipartition Property for Perplexity in Generative Language Models

### v3

@@ 2.3 A Modest Generalization
This theoretical analysis relies on Proposition 2.1 which only applies to toy models that generate strings of iid tokens. Of course, the simple AEP in Proposition 2.1 has been extended in various ways in the literature such as the following version (see (Tal et al., 2017), Appendix B.4), which we will make use of in the sequel.
@@
$\lim_{N\rightarrow\infty}\Pr[|l_{A^{\prime}}({\mathbf{A}}^{\prime}_{N})-H_{A^{\prime}}(N)|>\epsilon]\equiv 0,~{}\forall\epsilon>0,$
$\mathrm{where}~{}H_{A^{\prime}}(N)\equiv\frac{1}{N}\sum_{n=1}^{N}H(\alpha_{n})$ (10)
The regularity condition that the entropies of $\alpha_{n}(y)$ are uniformly bounded is trivially true if the token dictionary $\mathcal{T}$ is a finite set: $H(p(y))\leq\log_{2}|\mathcal{T}|$ for any distribution $p(y)$ over $y\in\mathcal{T}$.
@@
3 An Equipartition Property for Perplexity
We now propose a generalization of the theory described in Section 2.2 to a more interesting class of models. Consider a language model M and a random infinitely long text string ${\mathbf{Y}}$ generated by M, whose probabilities for a given prompt ${\mathbf{X}}$ is described by (2). Specifically, given a model M and a string ${\mathbf{Y}}$, (2) defines a sequence of probability distributions $p_{n}(y)$ over the set of tokens $y\in\mathcal{T}$. We will assume that the prompt ${\mathbf{X}}$ is fixed and omit it from our notation in the sequel.
@@ 3 An Equipartition Property for Perplexity
where $H(p)\equiv-\sum_{y\in\mathcal{T}}p(y)\log_{2}(p(y))$ is the entropy of the distribution $p(y)$ over $y\in\mathcal{T}$.
Note that the empirical entropy $h_{M}({\mathbf{Y}}_{N})$ is not the same as $\frac{1}{N}H\left(P({\mathbf{Y}}_{N})\right)$; the former expression is a random variable, whereas the latter is the entropy of $P({\mathbf{Y}}_{N})$ over ${\mathbf{Y}}_{N}\in\mathcal{T}^{N}$ and thus is a constant number. Indeed, we can show that
$\operatorname*{{E}}\left[l_{M}({\mathbf{Y}}_{N})\right]$ $\equiv\operatorname*{{E}}\left[h_{M}({\mathbf{Y}}_{N})\right]\equiv\frac{1}{N}H\left(P({\mathbf{Y}}_{N})\right)$ (12)
where the expectation is over random text strings ${\mathbf{Y}}_{N}$.
@@
Let the log-deviation $\lambda(p)$ of a probability distribution $p(y)$ over $y\in\mathcal{T}$ be defined as:
$\lambda(p)$ $\doteq\sqrt{S(p)-(H(p))^{2}},$
$\mathrm{where}~{}S(p)$ $\doteq\sum_{y\in\mathcal{T}}p(y)\left(\log_{2}p(y)\right)^{2}$ (13)
@@
We can again interpret $H(p_{n}),~{}\lambda(p_{n})$ as the conditional mean and standard deviation of $-\log_{2}(\Pr(Y_{n}))$ conditioned on ${\mathbf{Y}}_{N-1}$. Clearly, if the log-deviations $\lambda(p_{n})$ of the distributions $p_{n}(y)$ for a string ${\mathbf{Y}}$ are uniformly upper-bounded, $\lambda_{M}({\mathbf{Y}}_{N})$ asymptotically vanishes. We are now ready to state our main result.
@@
AEP for Perplexity. For a long text string ${\mathbf{Y}}$ i.e. $N\gg 1$ generated from a language model with a given prompt ${\mathbf{X}}$, with high probability, the log perplexity $l_{M}({\mathbf{Y}})$ is close to the empirical entropy $h_{M}({\mathbf{Y}})$. More precisely:
@@
We want to argue that $\lambda_{M}({\mathbf{Y}}_{N})$ is the standard deviation of $l_{M}({\mathbf{Y}}_{N})-h_{M}({\mathbf{Y}}_{N})$, and therefore asymptotically vanishing $\lambda_{M}({\mathbf{Y}}_{N})$ implies $l_{M}({\mathbf{Y}}_{N})\rightarrow h_{M}({\mathbf{Y}}_{N})$. This, however, is very wrong - for one thing $\lambda_{M}({\mathbf{Y}}_{N})$ is itself a random variable. In the sequel, we propose an Indifference Principle that allows us to carry through reasoning like the above to prove Proposition 3.1 and potentially other results. But first we present a direct proof of Proposition 3.1.
@@ 3.1 Formal Proof: an LLN for LLMs
We will use a version of the Weak Law of Large Numbers that applies to sequences of random variables with no restrictions on their dependence. Such a Law is presented in Appendix B (Proposition B.1). In fact, it turns out that a weaker version of Proposition B.1 namely Lemma B.2 is sufficient for our purposes.
@@
We note that the random variables $Z_{n}\doteq l_{n}$ satisfy all of the conditions for Lemma B.2, if we impose the restriction that they are strictly bounded i.e. $|l_{n}|<L$, for some some finite $L$. This is equivalent to requiring that all non-zero token probabilities $p_{n}(y)$ are above some lower bound e.g. $p_{n}(y)\geq 10^{-10}$. In practice, this is easily satisfied by any practical LLM. Then from Lemma B.2, we have for any $\epsilon>0$:
@@ 3.2 Informal Proof: Roads not Traveled
We now present an alternative proof of Proposition 3.1 using a more informal argument. Specifically, we introduce the conceptual device of an auxiliary language model that will allow us to apply the traditional Law of Large Numbers to a sequence of dependent random variables.
@@
Figure 1: Probability tree for the first 3 tokens of a language model M that selects each token $Y_{n}$ from a dictionary $\mathcal{T}$ of three tokens. Also shown in red is the probability tree of an auxiliary model $M^{\prime}_{\mathbf{S}}$ for a string ${\mathbf{S}}$ with $S_{1}=3,~{}S_{2}=1$ and so on.
3.2.1 Concept of an Auxiliary Language Model
@@
Let us consider a fixed string ${\mathbf{S}}$. For this particular string ${\mathbf{S}}$ and model M, we have a fixed set of distributions $q_{n}(y)\doteq\Pr(S_{n}=y|{\mathbf{S}}_{n-1})$. Now define an auxiliary generative model $M^{\prime}_{\mathbf{S}}$ that generates a random string ${\mathbf{Y}}^{\prime}\doteq[Y_{1}^{\prime},~{}Y_{2}^{\prime},~{}\dots,~{}Y_{N}^{\prime},~{}\dots]$ where the tokens $Y_{n}^{\prime}$ are generated independently from the distributions $q_{n}(y)$. If we represent a language model M as a probability tree, an auxiliary model $M^{\prime}_{\mathbf{S}}$ represents one specific subtree corresponding to the string ${\mathbf{S}}$ that defines the model. This is illustrated in Fig. 1. The auxiliary model $M^{\prime}_{\mathbf{S}}$ is much simpler than the general language model M, because its output tokens $Y^{\prime}_{n}$ are generated independently of other tokens. Our informal proof of Proposition 3.1 is based on the following idea.
@@
The Indifference Principle. For any random event $E$ defined entirely on the subtree of a particular string ${\mathbf{S}}$, $\Pr(E)$ will be the same for generative model M as for the auxiliary model $M^{\prime}_{\mathbf{S}}$.
The idea behind Proposition 3.2 is that while calculating $\Pr(E)$ for some event $E$, we can change any part of the probability tree that do not affect the event $E$ (“the roads not traveled”) to make our calculation easier e.g. by allowing us to work with an auxiliary language model $M^{\prime}_{\mathbf{S}}$ which generates strings of independent tokens. For our purposes, Proposition 3.2 applies to all events in the $\sigma$-algebra generated by the conditional likelihood random variables $l_{n}$ as defined in (16).
Informal Proof of Proposition 3.1..
Using the Proposition 3.2, we can evaluate the distribution of random variables $l_{n}$ assuming the tokens $Y_{n}$ are generated by the auxiliary model $M^{\prime}_{\mathbf{S}}$ instead of the original model M. Then the random variables $l_{m},~{}l_{n}$ are independent of each other and $\operatorname*{{E}}\left[l_{n}\right]\equiv H\left(q_{n}\right)$. The weak Law of Large Numbers for the sequence of independent zero-mean random variables $l_{n}-H\left(q_{n}\right)$ immediately gives us (17). ∎
Proposition 3.2 allows us to use traditional statistical laws that apply to sequences of independent random variables directly without having to prove bespoke versions of those laws such as Proposition B.1. For instance, we may be able to use Proposition 3.2 to formulate a version of the Central Limit Theorem for $\frac{1}{\lambda_{M}({\mathbf{S}}_{N})}\left(l_{M}({\mathbf{S}}_{N})-h_{M}({\mathbf{S}}_{N})\right)$. We delegate a full development of these ideas to future work and limit ourselves to one simple non-asymptotic result to illustrate the method.
3.3 Chebyshev Bound on Deviations
We note that for the auxiliary model $M^{\prime}_{\mathbf{S}}$, the variance of the conditional log likelihood random variable $l_{n}$ is $\lambda^{2}(q_{n})$ as defined in (13). Assuming independence of $l_{n}$ using Proposition 3.2, we can show that the variance of $l_{M}({\mathbf{S}}_{N})$ is $\lambda^{2}_{M}({\mathbf{S}}_{N})$ as defined in (14). We can then write using the Chebyshev’s Inequality (Cohen, 2015) for any $\alpha>0$:
$\Pr\left(\left|l_{M}({\mathbf{S}}_{N})-h_{M}({\mathbf{S}}_{N})\right|>\alpha\lambda_{M}({\mathbf{S}}_{N})\right)$ $\leq\frac{1}{\alpha^{2}}$ (18)
3.4 Typical Sets are Vanishingly Small
Let us now formally define the typical set $\mathcal{P}_{M}^{n}(\epsilon)\subset\mathcal{T}^{n}$ of typical strings ${\mathbf{Y}}_{n}$ for a model M with $\epsilon>0$:
@@ 3.4 Typical Sets are Vanishingly Small
$\mathcal{P}_{M}^{n}(\epsilon)$$\doteq\left\{{\mathbf{Y}}_{n}:\left|l_{M}({\mathbf{Y}}_{n})-h_{M}({\mathbf{Y}}_{n})\right|<\epsilon\right\}$ (19)
$\equiv\left\{{\mathbf{Y}}_{n}:\left|h_{M}({\mathbf{Y}}_{n})+\frac{1}{n}\log_{2}\left(\Pr({\mathbf{Y}}_{n})\right)\right|<\epsilon\right\}$
Proposition 3.1 asserts that with high probability a stochastic model M is constrained to only output typical strings from this set $\mathcal{P}_{M}^{n}(\epsilon)$ for any $\epsilon>0$ if we consider sufficiently long strings. This observation has profound consequences. We will now show one such consequence, namely that the typical set $\mathcal{P}_{M}^{n}(\epsilon)$ is very small:

### v4

@@ 2.2 A Modest Generalization
The simple AEP in Proposition 2.1 has been extended in various ways in the literature such as the following version (see Tal et al. (2017), Appendix B.4), which we will make use of in the sequel.
@@
$\lim_{N\rightarrow\infty}\Pr[|l_{A^{\prime}}({\mathbf{A}}^{\prime}_{N})-h_{A^{\prime}}(N)|>\epsilon]\equiv 0,~\forall\epsilon>0,\mathrm{where}~h_{A^{\prime}}(N)\equiv\frac{1}{N}\sum_{n=1}^{N}H(\alpha_{n})$ (9)
The regularity condition that the entropies of $\alpha_{n}(y)$ are uniformly bounded is trivially true if the token dictionary $\mathcal{Y}$ is a finite set: $H(p(y))\leq\log_{2}|\mathcal{Y}|$ for any distribution $p(y)$ over $y\in\mathcal{Y}$.
@@
3 An Un-Equipartition Property for Perplexity
We now propose a generalization of the theory described in Section 2.1 to a more interesting class of models. Consider a language model M and a random infinitely long text string $\mathbf{Y}$} generated by M, whose probabilities for a given prompt $\mathbf{X}$} is described by (2). Specifically, given a model M and a string $\mathbf{Y}$}, (2) defines a sequence of probability distributions $p_{n}(y)$ over the set of tokens $y\in\mathcal{Y}$. We will assume that the prompt $\mathbf{X}$} is fixed and omit it from our notation in the sequel.
@@ 3 An Un-Equipartition Property for Perplexity
where $H(p)\equiv-\sum_{y\in\mathcal{Y}}p(y)\log_{2}(p(y))$ is the entropy of the distribution $p(y)$ over $y\in\mathcal{Y}$.
Note that we make no assumptions about the existence of an “entropy rate” $\lim_{N\rightarrow\infty}h_{M}({\mathbf{Y}}_{N})$ for the model $M$. However, we can show that
$\operatorname*{{E}}\left[l_{M}({\mathbf{Y}}_{N})\right]$ $\equiv\operatorname*{{E}}\left[h_{M}({\mathbf{Y}}_{N})\right]\equiv\frac{1}{N}H\left(\Pr({\mathbf{Y}}_{N})\right)$ (11)
where $H\left(\Pr({\mathbf{Y}}_{N})\right)$ is the entropy of the random text strings ${\mathbf{Y}}_{N}$.
@@
Let the log-deviation $\lambda(p)$ of a probability distribution $p(y)$ over $y\in\mathcal{Y}$ be defined as:
$\lambda(p)$ $\doteq\sqrt{S(p)-(H(p))^{2}},~\mathrm{where}~S(p)\doteq\sum_{y\in\mathcal{Y}}p(y)\left(\log_{2}p(y)\right)^{2}$ (12)
@@
We can again interpret $H(p_{n}),~\lambda(p_{n})$ as the conditional mean and standard deviation of $\log_{2}(\Pr(Y_{n}))$ conditioned on ${\mathbf{Y}}_{N-1}$. Clearly, if the log-deviations $\lambda(p_{n})$ of the distributions $p_{n}(y)$ for a string $\mathbf{Y}$} are uniformly upper-bounded, $\lambda_{M}({\mathbf{Y}}_{N})$ asymptotically vanishes. We are now ready to state our main result.
@@
Un-Equipartition Property for Perplexity. For a long text string $\mathbf{Y}$} i.e. $N\gg 1$ generated from a language model with a given prompt $\mathbf{X}$}, with high probability, the log perplexity $l_{M}({\mathbf{Y}})$ is close to the empirical entropy $h_{M}({\mathbf{Y}})$. More precisely:
@@
We want to argue that $\lambda_{M}({\mathbf{Y}}_{N})$ is the standard deviation of $l_{M}({\mathbf{Y}}_{N})-h_{M}({\mathbf{Y}}_{N})$, and therefore asymptotically vanishing $\lambda_{M}({\mathbf{Y}}_{N})$ implies $l_{M}({\mathbf{Y}}_{N})\rightarrow h_{M}({\mathbf{Y}}_{N})$. This, however, is very wrong - for one thing $\lambda_{M}({\mathbf{Y}}_{N})$ is itself a random variable. In the sequel, we show an informal argument to carry through reasoning like the above. But first we present a direct proof of Proposition 3.1.
@@ 3.1 Formal Proof: an LLN for LLMs
We will use a version of the Weak Law of Large Numbers that applies to sequences of random variables with no restrictions on their dependence. Such a Law can be obtained using martingale difference ideas and an elementary derivation is presented in Appendix A.
@@
We note that the random variables $Z_{n}\doteq l_{n}$ satisfy all of the conditions for Lemma A.2, if we impose the restriction that they are strictly bounded i.e. $|l_{n}|<L$, for some some finite $L$. This is equivalent to requiring that all non-zero token probabilities $p_{n}(y)$ are above some lower bound e.g. $p_{n}(y)\geq 10^{-10}$. In practice, this is easily satisfied by any practical LLM. Then from Lemma A.2, we have for any $\epsilon>0$:
@@ 3.2 Informal Proof: Roads not Traveled
Figure 1: Probability tree for the first 3 tokens of a language model M that selects each token $Y_{n}$ from a dictionary $\mathcal{Y}$ of three tokens. Also shown in red is the probability tree of an auxiliary model $M^{\prime}_{\mathbf{S}}$ for a string $\mathbf{S}$} with $S_{1}=3,~S_{2}=1$ and so on.
@@
Consider a fixed string $\mathbf{S}$}. For this particular string $\mathbf{S}$} and model M, we have a fixed set of distributions $q_{n}(y)\doteq\Pr(S_{n}=y|{\mathbf{S}}_{n-1})$. Now define an auxiliary generative model $M^{\prime}_{\mathbf{S}}$ that generates a random string ${\mathbf{Y}}^{\prime}\doteq[Y_{1}^{\prime},~Y_{2}^{\prime},~\dots,~Y_{N}^{\prime},~\dots]$ where the tokens $Y_{n}^{\prime}$ are generated independently from the distributions $q_{n}(y)$. If we represent a language model M as a probability tree, an auxiliary model $M^{\prime}_{\mathbf{S}}$ represents one specific subtree corresponding to the string $\mathbf{S}$} that defines the model. This is illustrated in Fig. 1. The auxiliary model $M^{\prime}_{\mathbf{S}}$ is much simpler than the general language model M, because its output tokens $Y^{\prime}_{n}$ are generated independently of other tokens. Our informal proof of Proposition 3.1 is based on the following idea: while calculating $\Pr(E)$ for some event $E$, we can change any part of the probability tree that do not affect the event $E$ (“the roads not traveled”) to make our calculation easier e.g. by allowing us to work with an auxiliary language model $M^{\prime}_{\mathbf{S}}$ which generates strings of independent tokens.
Informal Proof of Proposition 3.1..
Assuming the tokens $Y_{n}$ are generated by the auxiliary model $M^{\prime}_{\mathbf{S}}$ (instead of the original model M), the random variables $l_{m},~l_{n}$ are independent and $\operatorname*{{E}}\left[l_{n}\right]\equiv H\left(q_{n}\right)$. Equation (16) then immediately follows from Proposition 2.2. ∎
This informal method allows us to treat the martingale property that the formal proof relies on as a quasi-independence condition to reason about the random variables $l_{n}$ intuitively. As an example, for the auxiliary model $M^{\prime}_{\mathbf{S}}$, from (12, 13), the variances of $l_{n},~l_{M}({\mathbf{S}}_{N})$ are respectively $\lambda^{2}(q_{n}),~\lambda^{2}_{M}({\mathbf{S}}_{N})$. We then have the following simple bound on deviations using the Chebyshev’s Inequality Cohen (2015):
$\Pr\left(\left|l_{M}({\mathbf{S}}_{N})-h_{M}({\mathbf{S}}_{N})\right|>\alpha\lambda_{M}({\mathbf{S}}_{N})\right)$ $\leq\frac{1}{\alpha^{2}},~\forall\alpha>0$ (17)
3.3 Typical Sets
We can now formally define the typical set $\mathcal{T}_{M}^{n}(\epsilon)\subset\mathcal{Y}^{n}$ of typical strings ${\mathbf{Y}}_{n}$ for a model M with $\epsilon>0$:
$\mathcal{T}_{M}^{n}(\epsilon)$ $\doteq\left\{{\mathbf{Y}}_{n}:\left|l_{M}({\mathbf{Y}}_{n})-h_{M}({\mathbf{Y}}_{n})\right|<\epsilon\right\}\equiv\left\{{\mathbf{Y}}_{n}:\left|h_{M}({\mathbf{Y}}_{n})+\frac{1}{n}\log_{2}\left(\Pr({\mathbf{Y}}_{n})\right)\right|<\epsilon\right\}$ (18)
Proposition 3.1 asserts that with high probability a stochastic model $M$ is constrained to only output typical strings from this set $\mathcal{T}_{M}^{n}(\epsilon)$ for any $\epsilon>0$ if we consider sufficiently long strings. In practical applications of the typical set concept, we almost invariably find that the typical set is vanishingly small. This is what makes the typical set concept so powerful: it identifies the small number of outcomes that actually matter, so we can focus our resources on them efficiently.
We will show that here too, the typical set must be vanishingly small under mild assumptions. Since the vast majority of token strings represent gibberish text, it is trivially true that any LLM that generates coherent text will not generate most token strings. Thus, the simple notion of smallness - that the typical set is a vanishing subset of all token strings - is not very interesting.
We want to show a much stronger result that even if we exclude gibberish sentences and only look at the set of grammatically correct texts, any practical LLM can only generate a vanishingly small fraction of all possible grammatical texts. This requires additional analytical machinery which we now introduce.
4 The Grammar Police
Let $\mathcal{G}(n)$ denote a dictionary of “grammatical” strings ${\mathbf{Y}}_{n}$ of length $n$ and $g(n)\doteq\frac{1}{n}\log_{2}|\mathcal{G}(n)|$. Since most token strings represent meaningless gibberish, we expect $|\mathcal{G}(n)|\ll|\mathcal{Y}|^{n}$ or $g(n)\ll\log_{2}|\mathcal{Y}|$ for any reasonable notion of grammar. However, we also want our dictionary to be rich enough to include sentences of unlimited length and complexity to avoid trivialities. Formally we require that $g(n)\geq g_{min}>0$. Beyond this, our analysis below is very general, which means we can define the notion of “grammatical strings” to be essentially anything we want.
4.1 Augmented Typical Sets for Language Models
We can now define the augmented typical set of the model $M$ for dictionary $\mathcal{G}(n)$ as:
$\mathcal{T}_{M,G}^{n}(\epsilon)$ $\doteq\mathcal{G}(n)\cap\mathcal{T}_{M}^{n}(\epsilon)\equiv\left\{{\mathbf{Y}}_{n}\in\mathcal{G}(n):\left|l_{M}({\mathbf{Y}}_{n})-h_{M}({\mathbf{Y}}_{n})\right|<\epsilon\right\}$ (19)
Let $p_{G}(n)\doteq\Pr({\mathbf{Y}}_{n}\in\mathcal{G}(n))$ be the probability that model $M$ generates a grammatically correct string ${\mathbf{Y}}_{n}\in\mathcal{G}(n)$. We will assume that $p_{G}(n)\geq p_{G}>0$ i.e. the model generates grammatical strings of any length with a non-zero probability. Informally, the model $M$ knows how to generate grammatically correct sentences when suitably prompted, but we do not require it to be perfect and it may still sometimes generate ungrammatical outputs. The probability that the model $M$ generates a grammatical string ${\mathbf{Y}}_{n}\in\mathcal{G}(n)$ can be written as:
$\Pr\big{(}{\mathbf{Y}}_{n}\big{|}{\mathbf{Y}}_{n}\in\mathcal{G}(n)\big{)}$ $=\frac{\Pr\big{(}{\mathbf{Y}}_{n}\big{)}}{p_{G}(n)}\equiv\frac{2^{-nl_{M}({\mathbf{Y}}_{n})}}{p_{G}(n)}$ (20)
We can use (20) to relate the perplexity of grammatical strings specifically to the entropy analogous to (11):
$\frac{1}{n}H\left(\Pr({\mathbf{Y}}_{n})\big{|}{\mathbf{Y}}_{n}\in\mathcal{G}(n)\right)$ $\equiv\frac{\log_{2}p_{G}(n)}{n}+\operatorname*{{E}}\left[l_{M}({\mathbf{Y}}_{n})\big{|}{\mathbf{Y}}_{n}\in\mathcal{G}(n)\right]\leq g(n)$ (21)
where the last inequality holds with equality for a model that selects all grammatical texts with equal probability: $\Pr({\mathbf{Y}}_{n})\equiv\frac{1}{|\mathcal{G}(n)|},~\forall{\mathbf{Y}}_{n}\in\mathcal{G}(n)$ (for which $p_{G}(n)=1$). Unlike (11), this does not tell us anything about the empirical entropy $h_{M}({\mathbf{Y}}_{n})$ of individual strings ${\mathbf{Y}}_{n}$. We will now show that we can safely ignore high entropy strings.
@@ 4.1 Augmented Typical Sets for Language Models
Entropies cannot be too large. The empirical entropy $h_{M}({\mathbf{Y}}_{n})$ of any long grammatical string ${\mathbf{Y}}_{n}\in\mathcal{G}(n)$ generated by the model $M$ is greater than $g(n)\equiv\frac{1}{n}\log_{2}|\mathcal{G}(n)|$ with vanishing probability:
$\lim_{n\rightarrow\infty}$ $\Pr\big{(}h_{M}({\mathbf{Y}}_{n})>g(n)+\epsilon\big{|}{\mathbf{Y}}_{n}\in\mathcal{G}(n)\big{)}=0,~\forall\epsilon>0$ (22)
Proof.
First we note that for sufficiently large $n$, we must have with high probability $\left|l_{M}({\mathbf{Y}}_{n})-h_{M}({\mathbf{Y}}_{n})\right|<\frac{\epsilon}{2}$. Then we have:
$\Pr\big{(}h_{M}({\mathbf{Y}}_{n})>g(n)+\epsilon\big{|}{\mathbf{Y}}_{n}\in\mathcal{G}(n)\big{)}$ $\leq\frac{1}{p_{G}(n)}\Pr\big{(}\mathcal{E}_{1}(n)\big{)}\leq\frac{1}{p_{G}}\Pr\big{(}\mathcal{E}_{1}(n)\big{)}$
@@
where $\mathcal{E}_{1}(n)\doteq\left\{{\mathbf{Y}}_{n}\in\mathcal{G}(n):l_{M}({\mathbf{Y}}_{n})>g(n)+\frac{\epsilon}{2}\right\}$. We can then bound this probability as:
$\Pr(\mathcal{E}_{1}(n))$ $=\sum_{{\mathbf{Y}}_{n}\in\mathcal{E}_{1}(n)}\Pr({\mathbf{Y}}_{n})\leq\sum_{{\mathbf{Y}}_{n}\in\mathcal{E}_{1}(n)}2^{-n(g(n)+\frac{\epsilon}{2})}\equiv\left|\mathcal{E}_{1}(n)\right|2^{-n(g(n)+\frac{\epsilon}{2})}<\left|\mathcal{G}(n)\right|2^{-n(g(n)+\frac{\epsilon}{2})}\equiv 2^{-\frac{n\epsilon}{2}}$
Thus we have shown that $\Pr\big{(}h_{M}({\mathbf{Y}}_{n})>g(n)+\epsilon\big{|}{\mathbf{Y}}_{n}\in\mathcal{G}(n)\big{)}<\frac{1}{p_{G}}2^{-\frac{n\epsilon}{2}}\rightarrow 0$ as $n\rightarrow\infty$. ∎
Proposition 4.1 allows us to purge high entropy strings out of the typical set while retaining all of its probability mass. Specifically, we define the purged typical set $\mathcal{P}_{M,G}^{n}(\epsilon)$ as:
$\mathcal{P}_{M,G}^{n}(\epsilon)$ $\doteq\mathcal{G}(n)\cap\mathcal{T}_{M}^{n}(\epsilon)\cap\mathcal{E}_{1}^{c}(n),~\mathrm{where}~\lim_{n\rightarrow\infty}\Pr\left(\mathcal{P}_{M,G}^{n}(\epsilon)|{\mathbf{Y}}_{n}\in\mathcal{G}(n)\right)=1,~\forall\epsilon>0$ (23)
4.2 The Essential Smallness of Typical Sets
For an entropy-maximizing model that chooses all grammatical strings with equal probability, the typical set includes all grammatical strings $\mathcal{P}_{M,G}^{n}(\epsilon)\equiv\mathcal{G}(n)$. We now formalize the assumption that practical LLMs do not do this, and for all such models, we will show that the typical set is a vanishingly small subset of $\mathcal{G}(n)$.
Assumption 4.1.
Practical LLMs are not entropy-maximizing. The empirical entropy $h_{M}({\mathbf{Y}}_{n})$ is strictly smaller than $g(n)$ for at least some fraction of grammatical typical strings. Specifically, there is an entropy gap $\Delta g>0$ such that $h_{M}({\mathbf{Y}}_{n})\leq g(n)-\Delta g$ for at least some non-zero fraction $\rho>0$ of typical strings ${\mathbf{Y}}_{n}$:
$\left|\mathcal{E}_{2}(n)\right|\geq\rho\left|\mathcal{P}_{M,G}^{n}(\epsilon)\right|,~\mathrm{where}~\mathcal{E}_{2}(n)\doteq\left\{{\mathbf{Y}}_{n}\in\mathcal{P}_{M,G}^{n}(\epsilon):h_{M}({\mathbf{Y}}_{n})\leq g(n)-\Delta g\right\}$ (24)